# Aligning Latent Spaces with Flow Priors

### v1

@@ 1 Introduction
Latent models like autoencoders (AEs) are a cornerstone of modern machine learning hinton2006reducing ; baldi2012autoencoders ; li2023comprehensive ; chen2023auto ; mienye2025deep . These models typically map high-dimensional observations to a lower-dimensional latent space, aiming to capture salient features and dependencies liou2014autoencoder ; meng2017relational . A highly desirable property of latent models is that the latent space should have structural properties, such as being close to a predefined target distribution salah2011contractive ; kingma2013auto ; yao2025reconstruction ; chen2024softvq . Such structure can incorporate domain-specific prior knowledge Khemakhem2019VariationalAA ; Raissi2019PhysicsinformedNN , enhance the interpretability of the latent spaceHiggins2016betaVAELB ; Chen2016InfoGANIR ; Kim2018DisentanglingBF , and facilitate latent space generation rombach2022high ; li2024autoregressive ; leng2025repa ; wen2025principal ; yu2024representation . While significant progress has been made, ensuring that the learned latent representations possess such desired structure remains a crucial challenge.
Traditional approaches to enforcing distributional conformity often involve minimizing divergences like the Kullback-Leibler (KL) divergence kingma2013auto ; rombach2022high . However, KL can be restrictive, particularly when the target prior is only implicitly defined (e.g., by samples). In latent generative modeling, the latent space is usually regularized with known prior distributions, such as the Gaussian distribution for Variational Autoencoders (VAE) kingma2013auto ; Esser2020TamingTF , and the categorical distribution for Vector Quantized VAE (VQ-VAE) van2017neural . Recent works qiu2025robust ; li2024imagefolder ; li2024xq ; chen2025masked ; yao2025reconstruction ; chen2024softvq have proposed to use pre-trained feature extractors as target distribution and directly optimize the latent distances, which are shown to be effective but computationally expensive and require per-sample features.
Recent advances in flow-based generative models lipman2022flow ; liu2022flow offer a promising avenue to capture complex target distributions. In this work, we address the question: Can we efficiently align a learnable latent space to an arbitrary target distribution using a pre-trained flow model as a prior? We answer this question affirmatively by proposing a novel framework that leverages a pre-trained flow model to define a computationally tractable alignment loss, which effectively guides the latents towards the target distribution.
@@
Our framework offers three key advantages. First, our approach enables alignment to arbitrary target distributions, even those implicitly defined by samples, overcoming the limitations of conventional methods that require explicit parametric priors. Second, the alignment loss acts as a direct surrogate for the log-likelihood of latents under the target distribution, providing a theoretically grounded objective that avoids heuristic metrics like cosine similarity used in per-sample feature matching chen2025masked ; yao2025reconstruction ; chen2024softvq . Third, our framework is computationally lightweight, requiring only a single forward pass through the pre-trained flow model during training, thereby bypassing the need for expensive adversarial optimization goodfellow2020generative , likelihood evaluations, or per-sample feature extraction qiu2025robust ; li2024imagefolder ; li2024xq .
We empirically validate the efficacy of our proposed alignment strategy through a series of experiments. We start with illustrative experiments in a controlled toy setting using a mixture of Gaussians to confirm that our alignment loss landscape indeed serves as a proxy for the log-likelihood of the latents under the target distribution. Then we demonstrate the scalability of our approach by conducting large-scale image generation experiments on ImageNet deng2009imagenet with diverse target distributions. Detailed discussions and ablation studies are provided to underscore the robustness and effectiveness.
@@ 2.1 Flow-based Models
Flow-based generative models have emerged as a powerful class of generative models esser2024scaling ; flux2024 ; chen2025goku ; Kingma2018GlowGF ; Zhai2024NormalizingFA ; Zhao2024FlowTurboTR ; shin2025deeply . They were first introduced as CNFs chen2018neural ; grathwohl2018ffjord that learn an invertible mapping between a simple base distribution (e.g., Gaussian) and a complex data distribution. Early works like NICE Dinh2014NICENI and Real NVP Dinh2016DensityEU introduced additive and affine coupling layers to construct invertible neural networks. A notable recent development is Flow Matching (FM) lipman2022flow ; albergo2022building ; liu2022flow ; neklyudov2023action ; heitz2023iterative ; tong2023improving , which simplifies the training by regressing a vector field against a target field derived from pairs of samples, avoiding the need for simulating ODEs during training. In ICTM Zhang2024FlowPF , flow priors of generative models have been employed for MAP estimation to solve linear inverse problems. Our work leverages flow-based models to learn complex distributions as a prior for latent space alignment.
@@ 2.2 Latent Space Alignment
The alignment of latent spaces with predefined distributions is a crucial aspect of representation learning. In VAE kingma2013auto , the latent space is typically regularized to follow a standard Gaussian distribution. Several approaches have been proposed to use more flexible priors, such as hierarchical VAEs Snderby2016LadderVA ; Vahdat2020NVAEAD or VAEs with inverse autoregressive flow (IAF) priors Kingma2016ImprovedVI . Another line of work focuses on aligning latent spaces with features extracted from pre-trained models qiu2025robust ; li2024imagefolder ; li2024xq ; chen2025masked ; yao2025reconstruction ; chen2024softvq ; kim2025democratizing ; zha2024language . Our method differs by utilizing a pre-trained flow model to define an expressive target and a novel alignment loss, avoiding expensive likelihoods, adversarial training, or direct per-sample feature comparison.
@@ 3.1 Flow-Based Models
Flow matching techniques lipman2022flow ; liu2022flow train $\bm{v}_{\theta}$ by minimizing its difference from a target velocity field. This target field is often defined by constructing a probability path $p_{t}(\bm{x})$ that interpolates between $p_{\mathrm{init}}$ and $p_{\mathrm{data}}$. A common choice is a conditional path $\bm{x}_{t}(\bm{x}_{0},\bm{x}_{1})$ defined for pairs $(\bm{x}_{0},\bm{x}_{1})$ sampled from $p_{\mathrm{init}}\times p_{\mathrm{data}}$. For instance, Rectified Flow uses a linear interpolation: $\bm{x}_{t}(\bm{x}_{0},\bm{x}_{1})=(1-t)\bm{x}_{0}+t\bm{x}_{1}$. The target velocity field corresponding to this path is $\bm{u}_{t}(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{1})=\bm{x}_{1}-\bm{x}_{0}$. The neural network $\bm{v}_{\theta}$ is then trained to predict this target field by minimizing the flow matching loss:
@@ 3.2 Likelihood Estimation with Flow Priors
Let $p_{1}^{\bm{v}_{\theta}}(\bm{x}_{1})$ denote the probability density at $t=1$ induced by the flow model $\bm{v}_{\theta}$ evolving from $p_{\mathrm{init}}$. Using the instantaneous change of variables formula, the log-likelihood of a sample $\bm{x}_{1}$ under this model can be computed by chen2018neural ; grathwohl2018ffjord :
@@ 4.1 Pipeline
The use of linear projection is theoretically supported by the Johnson-Lindenstrauss (JL) lemma johnson1984extensions . The JL lemma states that for a set of $N$ points in $\mathbb{R}^{d_{2}}$, a random linear mapping can preserve all pairwise Euclidean distances within a multiplicative distortion factor. The linear projection is defined by a matrix $\bm{W}\in\mathbb{R}^{d_{1}\times d_{2}}$. Assuming the target features $\bm{x}$ are appropriately normalized, we initialize the projection matrix $\bm{W}$ by sampling its entries from $\mathcal{N}(0,1/d_{2})$. This scaling helps ensure that the components of $\bm{x}$ also have approximately unit variance if the components of $\bm{x}$ are uncorrelated, thereby preserving key statistical properties.
@@ 4.3 Relating the Alignment Loss to an ELBO on Log-Likelihood
We establish this result by constructing a specific variational lower bound on $\log p_{1}^{\bm{v}_{\theta}}(\bm{y})$. Variational lower bounds for log-likelihoods in continuous-time generative models can be constructed by introducing a proposal distribution for the latents that could generate $\bm{y}$. Consider a family of "proposal" paths lipman2022flow , which are straight lines interpolating from an initial point $\bm{x}_{0}\sim p_{\mathrm{init}}$ to the given point $\bm{y}$:
@@
A known variational lower bound on $\log p_{1}^{\bm{v}_{\theta}}(\bm{y})$chen2018neural ; grathwohl2018ffjord ; liu2022flow can be written a:
@@ 5 Experiments
Figure 3: Illustration with a Mixture of Gaussians distribution. (a) Aligned latent variables $\bm{y}$ (red triangles) concentrate in low negative log-likelihood (NLL) regions of $p_{\text{data}}$ (blue dots; heatmap shows $-\log p_{\text{data}}$). (b) Alignment loss $\mathcal{L}_{\text{align}}$ heatmap mirrors the NLL landscape of $p_{\text{data}}$, with $p_{\text{data}}$ samples in low-$\mathcal{L}_{\text{align}}$ areas. (c) $\mathcal{L}_{\text{align}}$ (blue solid) and $-\log p_{\text{data}}(\bm{y})$ (red dashed) decline simultaneously in training, showing $\mathcal{L}_{\text{align}}$ serves as a proxy for maximizing the log-likelihood of $\bm{y}$ under $p_{\text{data}}$.
@@ 5.1 Toy Examples
We present a toy example as an illustrative experiment in a 2D setting. The target distribution, denoted $p_{\text{data}}$, is configured as a mixture of five isotropic Gaussians. Following the methodology outlined in Sec. 4.1, we first train a flow model $\bm{v}_{\theta}$ to map a standard 2D Normal distribution $\mathcal{N}(\bm{0},\bm{I})$ to $p_{\text{data}}$. This flow model $\bm{v}_{\theta}$ is implemented by a multi-layer perceptron (MLP) incorporating adaptive layer normalization for time modulation Peebles2022ScalableDM . Upon completion of training, the parameters $\theta$ of this flow model are frozen. Subsequently, instead of a parameterized model $G_{\phi}$, we directly initialize a set of learnable 2D variables as $\bm{y}$ and optimize them by minimizing the alignment loss $\mathcal{L}_{\text{align}}(\bm{y};\theta)$.
The results are presented in Fig. 3. Fig. 3 (a) compares the target distribution $p_{\text{data}}$ (blue point samples) with the optimized variables $\bm{y}$ (red triangles). The background visualizes the negative log-likelihood (NLL) of $p_{\text{data}}$, which is computed analytically. It is evident that $\bm{y}$ successfully converges to the high log-likelihood regions of $p_{\text{data}}$. Fig. 3 (b) displays the landscape of the alignment loss $\mathcal{L}_{\text{align}}$, which is estimated numerically with $\bm{v}_{\theta}$. The landscape mirrors the NLL surface of $p_{\text{data}}$ depicted in (a). Samples drawn from $p_{\text{data}}$ (blue dots) are concentrated in regions where $\mathcal{L}_{\text{align}}$ is low, suggesting that $\mathcal{L}_{\text{align}}$ effectively captures the underlying structure of the target distribution. Fig. 3 (c) illustrates $\mathcal{L}_{\text{align}}$ (blue solid line) and the true NLL $-\log p_{\text{data}}(\bm{y})$ (red dashed line) during the training of $\bm{y}$. The alignment loss and the NLL exhibit a strong positive correlation, decreasing concomitantly throughout the training process. More detailed toy examples can be found in Appendix B.
@@ 5.2 Image Generation
Prior work has demonstrated that aligning the latent space of AEs with semantic encoders can enhance generative model performance chen2025masked ; yao2025reconstruction ; chen2024softvq ; qiu2025robust . To validate this observation and further showcase the efficacy of our proposed method, we conduct large-scale image generation experiments on the ImageNet-1K deng2009imagenet dataset at $256\times 256$ resolution.
@@
Our AE architecture employs two Vision Transformer (ViT)-Large Dosovitskiy2020AnII models, each with 391M parameters, serving as the latent encoder and decoder, respectively. The encoder maps input images to a latent space of 64 tokens, each with dimension 32, striking a balance between reconstruction quality and computational efficiency. We impose token-level alignment on the latents. The alignment loss on the latents is set to $\lambda=0.01$ by default. We also incorporate conventional reconstruction loss, perceptual loss, and adversarial loss on the pixel outputs Esser2020TamingTF ; rombach2022high .
For the target distribution $p_{\text{data}}$, we investigate four distinct variants: low-level visual features from a VAE kingma2013auto ; Esser2020TamingTF , continuous semantic visual features from DinoV2 oquab2023dinov2 , discrete visual codebook embeddings from LlamaGen VQ sun2024autoregressive ; van2017neural , and textual embeddings from Qwen bai2023qwen . Their feature dimensions are 32, 768, 8, 896, respectively. The flow-based prior is modeled by a 6-layer MLP with 1024 hidden units, trained for 1 million steps using the AdamW loshchilov2017decoupled optimizer to match Assumption 1. More details can be found in Appendix C.
@@
Analogous to the toy example, we aim to correlate the alignment loss $\mathcal{L}_{\text{align}}$ with the NLL of latents under the target distribution $p_{\text{data}}$. Since the NLL is intractable for implicitly defined distributions, we estimate the density using $k$-nearest neighbors. The probability density $p(\bm{y})$ at a point $\bm{y}$ is inversely proportional to the volume of the hypersphere enclosing its $k^{\text{th}}$ nearest neighbor among the target samples. Consequently, the NLL can be estimated as $-\log p_{\text{data}}(\bm{y})\propto\log r_{k}(\bm{y})$ where $r_{k}(\bm{y})$ is the Euclidean distance to the $k^{\text{th}}$ neighbor, and $D$ is the dimension. We use $\log r_{k}(\bm{y})$ as our proxy measure for the NLL. We first index the set of target distribution samples using Faiss douze2024faiss . During the training, we periodically sample 10k points from the latent space and measure the alignment quality by averaging the $\log r_{k}(\bm{y})$.
@@
After demonstrating effective latent space alignment, we investigated its impact on generative model performance. We evaluated both reconstruction and generation capabilities on ImageNet using the MAR-B li2024autoregressive architecture. For MAR-B, we incorporated qk-norm dehghani2023scaling and replaced the diffusion head with a flow head to ensure stable training. We choose flow-based MAR-B as it does not favor continuous Gaussian-like latent structure like Diffusion models song2020denoising ; dhariwal2021diffusion ; Karras2022ElucidatingTD ; Nichol2021ImprovedDD ; rombach2022high do. To ensure an “apple-to-apple” comparison, configurations and hardware remained identical across all experiments, with the only difference being the specific AE used for each alignment variant.
@@
The results are presented in Tab. 1. Reconstruction performance was measured by rFID Heusel2017GANsTB and PSNR on the ImageNet validation set. Generation performance was assessed using FID, IS salimans2016improved , Precision, and Recall on 50k generated samples and the validation set, both with and without classifier-free guidance (CFG) ho2022classifier . Our key findings are:
1) Alignment vs. Reconstruction Trade-off: Latent space alignment typically degrades reconstruction quality (rFID, PSNR) compared to vanilla AEs, as constraints reduce capacity. SoftVQchen2024softvq excels among aligned methods due to its sample-level alignment. 2) Alignment Enhances Generation: Structured latent spaces improve generative metrics (FID, IS), but complexity is not decisive. Simple features (text embeddings like Qwen) may match the performance of richer visual features (DinoV2). 3) Optimal prior selection is open: No consensus exists on optimal priors. Low-dimensional discrete features (LlamaGen VQ) underperform, while cross-modal alignment (Qwen text embeddings) demonstrates transferable structural benefits. More discussions can be found in Appendix D.

### v2

@@ Abstract
This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000.
@@ 1 Introduction
Latent models like autoencoders (AEs) are a cornerstone of modern machine learning Hinton and Salakhutdinov (2006); Baldi (2012); Li et al. (2023); Chen and Guo (2023); Mienye and Swart (2025). These models typically map high-dimensional observations to a lower-dimensional latent space, aiming to capture salient features and dependencies Liou et al. (2014); Meng et al. (2017). A highly desirable property of latent models is that the latent space should have structural properties, such as being close to a predefined target distribution Salah et al. (2011); Kingma et al. (2013); Yao et al. (2025); Chen et al. (2024). Such structure can incorporate domain-specific prior knowledge Khemakhem et al. (2019); Raissi et al. (2019), enhance the interpretability of the latent spaceHiggins et al. (2016); Chen et al. (2016); Kim and Mnih (2018), and facilitate latent space generation Rombach et al. (2022); Li et al. (2024a); Leng et al. (2025); Wen et al. (2025); Yu et al. (2024). While significant progress has been made, ensuring that the learned latent representations possess such desired structure remains a crucial challenge.
Traditional approaches to enforcing distributional conformity often involve minimizing divergences like the Kullback-Leibler (KL) divergence Kingma et al. (2013); Rombach et al. (2022). However, KL can be restrictive, particularly when the target prior is only implicitly defined (e.g., by samples). In latent generative modeling, the latent space is usually regularized with known prior distributions, such as the Gaussian distribution for Variational Autoencoders (VAE) Kingma et al. (2013); Esser et al. (2020), and the categorical distribution for Vector Quantized VAE (VQ-VAE) Van Den Oord et al. (2017). Recent works Qiu et al. (2025); Li et al. (2024b, c); Chen et al. (2025a); Yao et al. (2025); Chen et al. (2024) have proposed to use pre-trained feature extractors as target distribution and directly optimize the latent distances, which are shown to be effective but computationally expensive and require per-sample features.
Recent advances in flow-based generative models Lipman et al. (2022); Liu et al. (2022) offer a promising avenue to capture complex target distributions. In this work, we address the question: Can we efficiently align a learnable latent space to an arbitrary target distribution using a pre-trained flow model as a prior? We answer this question affirmatively by proposing a novel framework that leverages a pre-trained flow model to define a computationally tractable alignment loss, which effectively guides the latents towards the target distribution.
@@
Our framework offers three key advantages. First, our approach enables alignment to arbitrary target distributions, even those implicitly defined by samples, overcoming the limitations of conventional methods that require explicit parametric priors. Second, the alignment loss acts as a direct surrogate for the log-likelihood of latents under the target distribution, providing a theoretically grounded objective that avoids heuristic metrics like cosine similarity used in per-sample feature matching Chen et al. (2025a); Yao et al. (2025); Chen et al. (2024). Third, our framework is computationally lightweight, requiring only a single forward pass through the pre-trained flow model during training, thereby bypassing the need for expensive adversarial optimization Goodfellow et al. (2020), likelihood evaluations, or per-sample feature extraction Qiu et al. (2025); Li et al. (2024b, c).
We empirically validate the efficacy of our proposed alignment strategy through a series of experiments. We start with illustrative experiments in a controlled toy setting using a mixture of Gaussians to confirm that our alignment loss landscape indeed serves as a proxy for the log-likelihood of the latents under the target distribution. Then we demonstrate the scalability of our approach by conducting large-scale image generation experiments on ImageNet Deng et al. (2009) with diverse target distributions. Detailed discussions and ablation studies are provided to underscore the robustness and effectiveness.
@@ 2.1 Flow-based Models
Flow-based generative models have emerged as a powerful class of generative models Esser et al. (2024); Labs (2024); Chen et al. (2025b); Kingma and Dhariwal (2018); Zhai et al. (2024); Zhao et al. (2024); Shin et al. (2025). They were first introduced as CNFs Chen et al. (2018); Grathwohl et al. (2018) that learn an invertible mapping between a simple base distribution (e.g., Gaussian) and a complex data distribution. Early works like NICE Dinh et al. (2014) and Real NVP Dinh et al. (2016) introduced additive and affine coupling layers to construct invertible neural networks. A notable recent development is Flow Matching (FM) Lipman et al. (2022); Albergo and Vanden-Eijnden (2022); Liu et al. (2022); Neklyudov et al. (2023); Heitz et al. (2023); Tong et al. (2023), which simplifies the training by regressing a vector field against a target field derived from pairs of samples, avoiding the need for simulating ODEs during training. In ICTM Zhang et al. (2024), flow priors of generative models have been employed for MAP estimation to solve linear inverse problems. Our work leverages flow-based models to learn complex distributions as a prior for latent space alignment.
@@ 2.2 Latent Space Alignment
The alignment of latent spaces with predefined distributions is a crucial aspect of representation learning. In VAE Kingma et al. (2013), the latent space is typically regularized to follow a standard Gaussian distribution. Several approaches have been proposed to use more flexible priors, such as hierarchical VAEs Sønderby et al. (2016); Vahdat and Kautz (2020) or VAEs with inverse autoregressive flow (IAF) priors Kingma et al. (2016). Another line of work focuses on aligning latent spaces with features extracted from pre-trained models Qiu et al. (2025); Li et al. (2024b, c); Chen et al. (2025a); Yao et al. (2025); Chen et al. (2024); Kim et al. (2025); Zha et al. (2024). Our method differs by utilizing a pre-trained flow model to define an expressive target and a novel alignment loss, avoiding expensive likelihoods, adversarial training, or direct per-sample feature comparison.
@@ 3.1 Flow-Based Models
Flow matching techniques Lipman et al. (2022); Liu et al. (2022) train $\bm{v}_{\theta}$ by minimizing its difference from a target velocity field. This target field is often defined by constructing a probability path $p_{t}(\bm{x})$ that interpolates between $p_{\mathrm{init}}$ and $p_{\mathrm{data}}$. A common choice is a conditional path $\bm{x}_{t}(\bm{x}_{0},\bm{x}_{1})$ defined for pairs $(\bm{x}_{0},\bm{x}_{1})$ sampled from $p_{\mathrm{init}}\times p_{\mathrm{data}}$. For instance, Rectified Flow uses a linear interpolation: $\bm{x}_{t}(\bm{x}_{0},\bm{x}_{1})=(1-t)\bm{x}_{0}+t\bm{x}_{1}$. The target velocity field corresponding to this path is $\bm{u}_{t}(\bm{x}_{t}|\bm{x}_{0},\bm{x}_{1})=\bm{x}_{1}-\bm{x}_{0}$. The neural network $\bm{v}_{\theta}$ is then trained to predict this target field by minimizing the flow matching loss:
@@ 3.2 Likelihood Estimation with Flow Priors
Let $p_{1}^{\bm{v}_{\theta}}(\bm{x}_{1})$ denote the probability density at $t=1$ induced by the flow model $\bm{v}_{\theta}$ evolving from $p_{\mathrm{init}}$. Using the instantaneous change of variables formula, the log-likelihood of a sample $\bm{x}_{1}$ under this model can be computed by Chen et al. (2018); Grathwohl et al. (2018):
@@ 4.1 Pipeline
The use of linear projection is theoretically supported by the Johnson-Lindenstrauss (JL) lemma Johnson et al. (1984). The JL lemma states that for a set of $N$ points in $\mathbb{R}^{d_{2}}$, a random linear mapping can preserve all pairwise Euclidean distances within a multiplicative distortion factor. The linear projection is defined by a matrix $\bm{W}\in\mathbb{R}^{d_{1}\times d_{2}}$. Assuming the target features $\bm{x}$ are appropriately normalized, we initialize the projection matrix $\bm{W}$ by sampling its entries from $\mathcal{N}(0,1/d_{2})$. This scaling helps ensure that the components of $\bm{x}$ also have approximately unit variance if the components of $\bm{x}$ are uncorrelated, thereby preserving key statistical properties.
@@ 4.3 Relating the Alignment Loss to an ELBO on Log-Likelihood
We establish this result by constructing a specific variational lower bound on $\log p_{1}^{\bm{v}_{\theta}}(\bm{y})$. Variational lower bounds for log-likelihoods in continuous-time generative models can be constructed by introducing a proposal distribution for the latents that could generate $\bm{y}$. Consider a family of "proposal" paths Lipman et al. (2022), which are straight lines interpolating from an initial point $\bm{x}_{0}\sim p_{\mathrm{init}}$ to the given point $\bm{y}$:
@@
A known variational lower bound on $\log p_{1}^{\bm{v}_{\theta}}(\bm{y})$Chen et al. (2018); Grathwohl et al. (2018); Liu et al. (2022) can be written a:
@@ 5 Experiments
Figure 3: Illustration with a Mixture of Gaussians distribution. (a) Aligned latent variables $\bm{y}$ (red triangles) concentrate in low negative log-likelihood (NLL) regions of $p_{\text{data}}$ (blue dots; heatmap shows $\log p_{\text{data}}$). (b) Alignment loss $\mathcal{L}_{\text{align}}$ heatmap mirrors the NLL landscape of $p_{\text{data}}$, with $p_{\text{data}}$ samples in low-$\mathcal{L}_{\text{align}}$ areas. (c) $\mathcal{L}_{\text{align}}$ (blue solid) and $\log p_{\text{data}}(\bm{y})$ (red dashed) decline simultaneously in training, showing $\mathcal{L}_{\text{align}}$ serves as a proxy for maximizing the log-likelihood of $\bm{y}$ under $p_{\text{data}}$.
@@ 5.1 Toy Examples
We present a toy example as an illustrative experiment in a 2D setting. The target distribution, denoted $p_{\text{data}}$, is configured as a mixture of five isotropic Gaussians. Following the methodology outlined in Sec. 4.1, we first train a flow model $\bm{v}_{\theta}$ to map a standard 2D Normal distribution $\mathcal{N}(\bm{0},\bm{I})$ to $p_{\text{data}}$. This flow model $\bm{v}_{\theta}$ is implemented by a multi-layer perceptron (MLP) incorporating adaptive layer normalization for time modulation Peebles and Xie (2022). Upon completion of training, the parameters $\theta$ of this flow model are frozen. Subsequently, instead of a parameterized model $G_{\phi}$, we directly initialize a set of learnable 2D variables as $\bm{y}$ and optimize them by minimizing the alignment loss $\mathcal{L}_{\text{align}}(\bm{y};\theta)$.
The results are presented in Fig. 3. Fig. 3 (a) compares the target distribution $p_{\text{data}}$ (blue point samples) with the optimized variables $\bm{y}$ (red triangles). The background visualizes the negative log-likelihood (NLL) of $p_{\text{data}}$, which is computed analytically. It is evident that $\bm{y}$ successfully converges to the high log-likelihood regions of $p_{\text{data}}$. Fig. 3 (b) displays the landscape of the alignment loss $\mathcal{L}_{\text{align}}$, which is estimated numerically with $\bm{v}_{\theta}$. The landscape mirrors the NLL surface of $p_{\text{data}}$ depicted in (a). Samples drawn from $p_{\text{data}}$ (blue dots) are concentrated in regions where $\mathcal{L}_{\text{align}}$ is low, suggesting that $\mathcal{L}_{\text{align}}$ effectively captures the underlying structure of the target distribution. Fig. 3 (c) illustrates $\mathcal{L}_{\text{align}}$ (blue solid line) and the true NLL $\log p_{\text{data}}(\bm{y})$ (red dashed line) during the training of $\bm{y}$. The alignment loss and the NLL exhibit a strong positive correlation, decreasing concomitantly throughout the training process. More detailed toy examples can be found in Appendix B.
@@ 5.2 Image Generation
Prior work has demonstrated that aligning the latent space of AEs with semantic encoders can enhance generative model performance Chen et al. (2025a); Yao et al. (2025); Chen et al. (2024); Qiu et al. (2025). To validate this observation and further showcase the efficacy of our proposed method, we conduct large-scale image generation experiments on the ImageNet-1K Deng et al. (2009) dataset at $256\times 256$ resolution.
@@
Our AE architecture employs two Vision Transformer (ViT)-Large Dosovitskiy et al. (2020) models, each with 391M parameters, serving as the latent encoder and decoder, respectively. The encoder maps input images to a latent space of 64 tokens, each with dimension 32, striking a balance between reconstruction quality and computational efficiency. We impose token-level alignment on the latents. The alignment loss on the latents is set to $\lambda=0.01$ by default. We also incorporate conventional reconstruction loss, perceptual loss, and adversarial loss on the pixel outputs Esser et al. (2020); Rombach et al. (2022).
For the target distribution $p_{\text{data}}$, we investigate four distinct variants: low-level visual features from a VAE Kingma et al. (2013); Esser et al. (2020), continuous semantic visual features from DinoV2 Oquab et al. (2023), discrete visual codebook embeddings from LlamaGen VQ Sun et al. (2024); Van Den Oord et al. (2017), and textual embeddings from Qwen Bai et al. (2023). Their feature dimensions are 32, 768, 8, 896, respectively. The flow-based prior is modeled by a 6-layer MLP with 1024 hidden units, trained for 1 million steps using the AdamW Loshchilov and Hutter (2017) optimizer to match Assumption 1. More details can be found in Appendix C.
@@
Analogous to the toy example, we aim to correlate the alignment loss $\mathcal{L}_{\text{align}}$ with the NLL of latents under the target distribution $p_{\text{data}}$. Since the NLL is intractable for implicitly defined distributions, we estimate the density using $k$-nearest neighbors. The probability density $p(\bm{y})$ at a point $\bm{y}$ is inversely proportional to the volume of the hypersphere enclosing its $k^{\text{th}}$ nearest neighbor among the target samples. Consequently, the NLL can be estimated as $\log p_{\text{data}}(\bm{y})\propto\log r_{k}(\bm{y})$ where $r_{k}(\bm{y})$ is the Euclidean distance to the $k^{\text{th}}$ neighbor, and $D$ is the dimension. We use $\log r_{k}(\bm{y})$ as our proxy measure for the NLL. We first index the set of target distribution samples using Faiss Douze et al. (2024). During the training, we periodically sample 10k points from the latent space and measure the alignment quality by averaging the $\log r_{k}(\bm{y})$.
@@
After demonstrating effective latent space alignment, we investigated its impact on generative model performance. We evaluated both reconstruction and generation capabilities on ImageNet using the MAR-B Li et al. (2024a) architecture. For MAR-B, we incorporated qk-norm Dehghani et al. (2023) and replaced the diffusion head with a flow head to ensure stable training. We choose flow-based MAR-B as it does not favor continuous Gaussian-like latent structure like Diffusion models Song et al. (2020); Dhariwal and Nichol (2021); Karras et al. (2022); Nichol and Dhariwal (2021); Rombach et al. (2022) do. To ensure an “apple-to-apple” comparison, configurations and hardware remained identical across all experiments, with the only difference being the specific AE used for each alignment variant.
@@
The results are presented in Tab. 1. Reconstruction performance was measured by rFID Heusel et al. (2017) and PSNR on the ImageNet validation set. Generation performance was assessed using FID, IS Salimans et al. (2016), Precision, and Recall on 50k generated samples and the validation set, both with and without classifier-free guidance (CFG) Ho and Salimans (2022). Our key findings are:
1) Alignment vs. Reconstruction Trade-off: Latent space alignment typically degrades reconstruction quality (rFID, PSNR) compared to vanilla AEs, as constraints reduce capacity. SoftVQChen et al. (2024) excels among aligned methods due to its sample-level alignment. 2) Alignment Enhances Generation: Structured latent spaces improve generative metrics (FID, IS), but complexity is not decisive. Simple features (text embeddings like Qwen) may match the performance of richer visual features (DinoV2). 3) Optimal prior selection is open: No consensus exists on optimal priors. Low-dimensional discrete features (LlamaGen VQ) underperform, while cross-modal alignment (Qwen text embeddings) demonstrates transferable structural benefits. More discussions can be found in Appendix D.